### v1

@@ I Introduction
From the view of particle physics, one of the most attractive schemes to solve dark matter problem is WIMPs (weak interaction massive particles), where the stable dark matter particle is at a weak scale and the observed dark matter relic density is generated via the so-called Freeze-out mechanismChiu (1966). In addition, scenarios in which two or more different particles contribute to the dark matter density— multi-component dark matter model are also possible Belanger et al. (2022); Alguero et al. (2024); Bélanger et al. (2023), where the total relic density of two or more dark matter particles contributes to the experiment results. For the different constituents of dark matter in the multi-component dark matter model, the corresponding density can be generated not just via the Freeze-out mechanism but also the Freeze-in mechanism Hall et al. (2010). Correspondingly, the components of dark matter can be WIMPs, FIMPs( Feebly interaction massive particles), or both. Among the multi-component dark matter models, the scalar fields that are simultaneously stabilized by a single $Z_{N}$ symmetry as dark matter particles are particularly appealingDeshpande and Ma (1978); Ma (2006); Barbieri et al. (2006); Lopez Honorez et al. (2007); Bélanger et al. (2014); Belanger et al. (2012); Díaz Sáez and Ghorbani (2023), these complex scalars are SM singlets but have different charges under the discrete symmetry. This symmetry, in turn, could be a remnant of a spontaneously broken $U(1)$ gauge symmetry and thus be related to gauge extensions of the SM Martin (1992); Krauss and Wilczek (1989).
@@
As for the neutrino mass problem, one can introduce new particles to the SM to generate the tiny neutrino mass with the so-called seesaw mechanism. Generally speaking, we have three kinds of seesaw mechanisms: Type-I seesaw Minkowski (1977); Mohapatra and Senjanovic (1980) (introducing new right-handed neutrinos with Majorana mass terms), Type-II seesaw Magg and Wetterich (1980); Lazarides et al. (1981); Ashery et al. (1981); Ma and Sarkar (1998); Konetschny and Kummer (1977); Schechter and Valle (1980); Cheng and Li (1980); Bilenky et al. (1980)(introducing new triplet scalars) and Type-III seesaw Foot et al. (1989); Ma (1998); Ma and Roy (2002)(introducing new fermion triplets). For the Type-II seesaw mechanism, we have a new triplet scalar $\Delta\equiv(\delta^{++},\delta^{+},\delta^{0})$ with a non-zero vacuum expectation value ($v_{\Delta}$). Since the triplet couples to SM gauge bosons and leptons directly, rich LHC phenomenology for a low seesaw scale can be found in the Type-II seesaw Akeroyd et al. (2008); Fileviez Perez et al. (2008a); del Aguila and Aguilar-Saavedra (2009); Melfo et al. (2012). What’s more, the light triplet can modify the Higgs-to-Z+photon Djouadi et al. (1998); Carena et al. (2012); Chen et al. (2013a) and Higgs-to-diphoton Melfo et al. (2012); Arhrib et al. (2012); Akeroyd and Moretti (2012) decay rates, which are correlated in the type-II seesaw model for most of the viable parameter space Bhupal Dev et al. (2013); Chen et al. (2013b).
@@
In this work, we consider the $Z_{5}$ two-component dark matter in the Type-II seesaw mechanism. The case of $Z_{5}$ two-component dark matter in the SM has been discussed in Bélanger et al. (2020); Qi and Sun (2023a), where two singlet scalars carrying $Z_{5}$ charge are introduced to the SM with the Higgs-portal couplings as dark matter, and we have new trilinear and quartic couplings of the two scalars due to the $Z_{5}$ symmetry, which can induce annihilation, semi-annihilation and conversion processes related to dark matter density. One feature of such a model is that the entire range of dark matter masses is allowed, and the dark matter density is always dominated by the lighter component. In this work, we have new couplings of dark matter with the triplet scalar so that new annihilation processes of the two singlet scalars are possible, which means the lighter component does not necessarily determine dark matter density. On the other hand, when dark matter is embedded in the Type-II seesaw, its interactions with the triplet scalar field $\Delta$ which dominantly decays to leptonic final states for a small triplet vacuum expectation value $v_{\Delta}<0.1$ MeV can contribute to fit the excess of positron-electron flux according to the AMS-02 data for the leptophilic nature of dark matter. Concretely speaking, the introduced triplet in the type-II seesaw mechanism can play an important role in exploring the observed excess of cosmic rays observed in the electron-positron flux measured by the AMS-02 Aguilar et al. (2014), Fermi-LAT Abdollahi et al. (2017), and DAMPE Ambrosi et al. (2017)experiments, arising from the leptonic decays of such a triplet during DM annihilation Dev et al. (2014); Li et al. (2018a, b). The W and Z boson pair from s-channel DM annihilation with subsequent decay may lead to an inappropriate antiproton spectrum measured by AMS and PAMELA in cosmic rays, which constrains the coupling between SM Higgs and dark matter. Note that the boost factor (BF) is also necessary to consider, which may come from large inhomogeneities in the dark matter distribution or due to the so-called Breit-Wigner enhancement mechanism in particle physics Ibe et al. (2009); March-Russell and West (2009); Guo and Wu (2009). Related discussion about cosmic ray excess in the Type-II seesaw mechanism can be found in Dev et al. (2014); Qi and Sun (2023b). In this paper, we consider the two-component dark matter case in the Type-II seesaw mechanism where both components of the dark matter can contribute to the possible positron-electron excess. As we mentioned above, either the light or the heavy component can be dominant in the model, and we will have different results to fit the observed positron-electron excess in both cases.
@@
The paper is arranged as follows, in Sec. II, we give the model framework of $Z_{5}$ two-component dark matter in the Type-II seesaw mechanism. In Sec. III, we discuss the dark matter phenomenology in the model. In Sec. IV, we consider the anti-proton spectrum and electron/positron excess in our framework, and we give a summary in the last part.
@@ II Model description
where $\phi_{1}$ and $\phi_{2}$ are dark matter carrying $Z_{5}$ charge, and we have the transform of $SM\to SM$, $\phi_{1}\to e^{i2\pi/5}\phi_{1}$ and $\phi_{2}\to e^{i4\pi/5}\phi_{2}$ under the $Z_{5}$ symmetry. The new trilinear and quartic couplings denoted $\mu_{si}$ and $\lambda_{3i}~{}({i=1,2})$ respectively.
@@
III Dark matter phenomenology
@@ III.2 Parameter dependence
III.3 The viable parameter space
@@
IV Anti-proton spectrum and electron/positron excess
Figure 12: Background and background + DM (Bkg + DM) of the cosmic-ray antiproton flux with the four benchmark values respectively, where different colored lines correspond to the results of BF taking different values. The data points in both pictures label the PAMELA Adriani et al. (2010) and AMS Aguilar et al. (2019) measurements.
Figure 13: comparison of the positron fraction observed by AMS-02 Aguilar et al. (2014) in the IH scenario (the purple lines) and NH scenario (the blue lines) with the four benchmark values and fixed BF value. The data points stand for the AMS-02, and the green lines represent the Background.
@@ IV Anti-proton spectrum and electron/positron excess
Table 2: Benchmark parameters satisfying relic density constraint and direct detection constraint, where Case I (Case III) and Case II (Case IV) correspond to the light (heavy) component dominant in dark matter relic density.
@@
We show the results of the background and Bkg + DM for $\Phi_{\bar{p}}$ in Fig. 12 with the above four cases respectively, where the PAMELA Adriani et al. (2010) results impose strong restrictions on the BF. In Fig. 12(a) and Fig. 12(b), we give the results of the light component dominant in dark matter relic density, while in Fig. 12(c) and Fig. 12(d), we give the heavy component cases. The allowed maximum values of BF are about 2000,400 and 2000 respectively according to pictures (a),(b) and (d) of Fig. 12. For the result of Case III (corresponding to Fig. 12(c)), since the couplings $\lambda_{s1h}$ and $\lambda_{s2h}$ are so small that the contribution of DM particles annihilating to W and Z boson pairs through the s-channel Higgs exchange can be negligible, therefore, the allowed BF value is more flexible compared with other cases.
@@
In Fig. 13, we give the corresponding results of the comparison of the positron fraction observed by the AMS-02 with the four benchmark values, where the positron fraction is defined by $\Phi_{e^{+}}/(\Phi_{e^{-}}+\Phi_{e^{+}})$. In the model, the final states of DM particle annihilation will depend on whether the neutrino mass is the inverted hierarchy (IH) or the normal hierarchy (NH). The annihilation production is mainly electron final states for the IH scenario, but tau final states for the NH scenario instead. In Fig. 13, the blue lines represent the results of the NH scenario and the purple lines correspond to the IH scenario. When the triplets decay mainly into electron/positron final states, the positron flux arising from DM annihilation will rise sharply above the background as long as we go to higher energies and will eventually drop to the background level, while for the NH scenario, the triplets mostly decay into taus and subsequently decay to electrons, and one can obtain a much softer energy spectrum of the positron flux before eventually dropping to the background level as we can see in Fig. 13(a) and Fig. 13(c). In Fig. 13(b) and Fig. 13(d) corresponding to Case II and Case IV, the results are above the background (without DM) but much lower than the AMS-02 data. The reason is that in those two cases, $m_{1}<M_{\Delta}<m_{2}$, and the annihilation process of $\phi_{1}+\phi_{1}^{\dagger}\to\Delta\Delta$ is kinetically suppressed, while the coupling $\lambda_{s2d}$ related to $\phi_{2}+\phi_{2}^{\dagger}\to\Delta\Delta$ is so small that contribution of dark matter annihilation to triplet is also suppressed. On the other hand, the NH scenario can give a better fit with the AMS-02 compared with the IH scenario as we can see in Fig. 13(a) and Fig. 13(c), which correspond to light component dominant and heavy component dominant in the model respectively. As we mentioned above, the positron energy spectrum is harder in the IH scenario since it comes directly from the $\Delta$ decay, while the spectrum is softer in the NH case, because it comes from muon and tau decays. The rise in the AMS-02 positron fraction becomes softer toward higher energies, therefore, it is easier to fit both low- and high-energy bins for the NH case compared to the IH case. In our model, both light components dominant in dark matter relic density and the heavy component case can give the proper results as we can see in Fig. 13(a) and Fig. 13(c) with $M_{\Delta}<m_{1}<m_{2}$, which is consistent with the results of dark matter in the Type-II seesaw mechanism.
V summary and outlook

### v5

@@ I Introduction
From the view of particle physics, one of the most attractive schemes to solve dark matter problem is WIMPs (weak interaction massive particles), where the stable dark matter particle is at a weak scale and the observed dark matter relic density is generated via the so-called Freeze-out mechanismChiu (1966). In addition, scenarios in which two or more different particles contribute to the dark matter density— multi-component dark matter model are also possible Belanger et al. (2022); Alguero et al. (2024); Bélanger et al. (2023); Bhattacharya et al. (2017, 2020, 2023, 2022), where the total relic density of two or more dark matter particles contributes to the experiment results. For the different constituents of dark matter in the multi-component dark matter model, the corresponding density can be generated not just via the Freeze-out mechanism but also the Freeze-in mechanism Hall et al. (2010). Correspondingly, the components of dark matter can be WIMPs, FIMPs( Feebly interaction massive particles), or both. Among the multi-component dark matter models, the scalar fields that are simultaneously stabilized by a single $Z_{N}$ symmetry as dark matter particles are particularly appealingDeshpande and Ma (1978); Ma (2006); Barbieri et al. (2006); Lopez Honorez et al. (2007); Bélanger et al. (2014); Belanger et al. (2012); Díaz Sáez and Ghorbani (2023), these complex scalars are SM singlets but have different charges under the discrete symmetry. This symmetry, in turn, could be a remnant of a spontaneously broken $U(1)$ gauge symmetry and thus be related to gauge extensions of the SM Martin (1992); Krauss and Wilczek (1989).
@@
As for the neutrino mass problem, one can introduce new particles to the SM to generate the tiny neutrino mass with the so-called seesaw mechanismSchechter and Valle (1980); Schechter and Valle (1982a, b). Generally speaking, we have three kinds of seesaw mechanisms: Type-I seesaw Minkowski (1977); Mohapatra and Senjanovic (1980) (introducing new right-handed neutrinos with Majorana mass terms), Type-II seesaw Magg and Wetterich (1980); Lazarides et al. (1981); Ashery et al. (1981); Ma and Sarkar (1998); Konetschny and Kummer (1977); Schechter and Valle (1980); Cheng and Li (1980); Bilenky et al. (1980)(introducing new triplet scalars) and Type-III seesaw Foot et al. (1989); Ma (1998); Ma and Roy (2002)(introducing new fermion triplets). For the Type-II seesaw mechanism, we have a new triplet scalar $\Delta\equiv(\delta^{++},\delta^{+},\delta^{0})$ with a non-zero vacuum expectation value ($v_{\Delta}$). Since the triplet couples to SM gauge bosons and leptons directly, rich LHC phenomenology for a low seesaw scale can be found in the Type-II seesaw Akeroyd et al. (2008); Fileviez Perez et al. (2008a); del Aguila and Aguilar-Saavedra (2009); Melfo et al. (2012). What’s more, the light triplet can modify the Higgs-to-Z+photon Djouadi et al. (1998); Carena et al. (2012); Chen et al. (2013a) and Higgs-to-diphoton Melfo et al. (2012); Arhrib et al. (2012); Akeroyd and Moretti (2012) decay rates, which are correlated in the type-II seesaw model for most of the viable parameter space Bhupal Dev et al. (2013); Chen et al. (2013b).
@@
In this work, we consider the $Z_{5}$ two-component dark matter in the Type-II seesaw mechanism. The case of $Z_{5}$ two-component dark matter in the SM has been discussed in Bélanger et al. (2020); Qi and Sun (2023a), where two singlet scalars carrying $Z_{5}$ charge are introduced to the SM with the Higgs-portal couplings as dark matter, and we have new trilinear and quartic couplings of the two scalars due to the $Z_{5}$ symmetry, which can induce annihilation, semi-annihilation and conversion processes related to dark matter density. One feature of such a model is that the entire range of dark matter masses is allowed, and the dark matter density is always dominated by the lighter component. In this work, we have new couplings of dark matter with the triplet scalar so that new annihilation processes of the two singlet scalars are possible, which means the lighter component does not necessarily determine dark matter density. On the other hand, when dark matter is embedded in the Type-II seesaw, its interactions with the triplet scalar field $\Delta$ which dominantly decays to leptonic final states for a small triplet vacuum expectation value $v_{\Delta}<0.1$ MeV can contribute to fit the excess of positron-electron flux according to the AMS-02 data for the leptophilic nature of dark matter. Concretely speaking, the introduced triplet in the type-II seesaw mechanism can play an important role in exploring the observed excess of cosmic rays observed in the electron-positron flux measured by the AMS-02 Aguilar et al. (2014), Fermi-LAT Abdollahi et al. (2017), and DAMPE Ambrosi et al. (2017)experiments, arising from the leptonic decays of such a triplet during DM annihilation Dev et al. (2014); Li et al. (2018a, b). The W and Z boson pair from s-channel DM annihilation with subsequent decay may lead to an inappropriate antiproton spectrum measured by AMS and PAMELA in cosmic rays, which constrains the coupling between SM Higgs and dark matter. Note that this model requires a larger boost factor (BF) of order $10^{3}-10^{4}$ in the DM annihilation rate to explain the observed positron excess Dev et al. (2014), which may come from small-scale inhomogeneities in the dark matter distribution which cannot be excluded even with the highest resolution numerical simulations available at presentKuhlen et al. (2012),or due to the so-called Breit-Wigner enhancement mechanism in particle physics Ibe et al. (2009); March-Russell and West (2009); Guo and Wu (2009). Related discussion about cosmic ray excess in the Type-II seesaw mechanism can be found in Dev et al. (2014); Qi and Sun (2023b). In this paper, we consider the two-component dark matter case in the Type-II seesaw mechanism where both components of the dark matter can contribute to the possible positron-electron excess. As we mentioned above, either the light or the heavy component can be dominant in the model, and we will have different results to fit the observed positron-electron excess in both cases.
@@
The paper is arranged as follows, in Sec. II, we give the model framework of $Z_{5}$ two-component dark matter in the Type-II seesaw mechanism. In Sec. IV, we discuss the dark matter phenomenology in the model. In Sec. V, we consider the anti-proton spectrum and electron/positron excess in our framework, and we give a summary in the last part.
@@ II Model description
where $\phi_{1}$ and $\phi_{2}$ are dark matter carrying $Z_{5}$ charge with $X_{\phi_{1}}=1/5$ and $X_{\phi_{2}}=2/5$, and we have the transform of $SM\to SM$, $\phi_{1}\to e^{i2\pi/5}\phi_{1}$ and $\phi_{2}\to e^{i4\pi/5}\phi_{2}$ under the $Z_{5}$ symmetry. The new trilinear and quartic couplings denoted $\mu_{si}$ and $\lambda_{3i}~{}({i=1,2})$ respectively.
@@
III Theoritical constraints
In this part, we consider the theoritical constraints on the parameter space including perturbativity, perturbative unitarity and vacuum stability.
III.1 Perturbativity
To illustrate the theoretical bounds from the perturbativity behavior of the dimensionless scalar quartic couplings, we follow the definitions in Refs. Bélanger et al. (2014); Lerner and McDonald (2009). As to the case of an unrotated basis, the vertices from the potential must be less than $4\pi$ to make sure that the tree-level contributions are larger than the one-loop level quantum corrections. This condition will give the constraints on the quartic couplings in the potential, which are
$|\lambda_{1}+\lambda_{4}|<4\pi,|\lambda_{1}+\frac{\lambda_{4}}{2}|<4\pi,|6(\lambda_{2}+\lambda_{3})|<4\pi,|2\lambda_{2}|<4\pi,|\sqrt{2}\lambda_{3}|<4\pi,|\lambda_{sid}|<4\pi,$
$|\lambda_{sih}|<4\pi|,|4\lambda_{4i}|<4\pi,|3\lambda_{3i}|<4\pi,|\lambda_{412}|<4\pi$ (24)
where $i=1,2$.
III.2 Perturbative unitarity
The tree-level unitarity from two-body scalar-scalar scattering processes gives another bound on the quartic couplings in the potential. When the collision energy $\sqrt{s}$ becomes larger, the processes will be dominated by the terms of quartic contact interaction. Although the trilinear couplings that are contributed to scattering should be included at finite collision energy Hektor et al. (2019); Schuessler and Zeppenfeld (2007), for simplicity, we only calculate the unitarity constraints with the following scenario: $s\to+\infty$. The s-wave scattering amplitudes lie in the perturbative unitarity limit, giving the constraint of the scalar-scalar scattering S-matrix values: $|\mathrm{Re}\mathcal{M}_{i}|\leqslant\frac{1}{2}$. The perturbative unitarity in the type-II seesaw model has been studied by decomposing the matrix S by the mutually unmixed sets of channels with definite charge and CP states Arhrib et al. (2011). We extend the way of decomposing by considering the $Z_{5}$ symmetry with $X_{\phi_{1}}=1/5$ singlet $\phi_{1}$ and $X_{\phi_{2}}=2/5$ singlet $\phi_{2}$ introduced in our model. The matrix S can be decomposed into seven submatrix blocks structured in terms of electric charges and $Z_{5}$ charges in the initial/final states.
In Appendix. A, we display the initial/final states $E_{i}$ and the corresponding scattering submatrix $\mathcal{M}_{i}$. The corresponding eigenvalues $e^{j}_{i}$ of each submatrix are then calculated. The limit from perturbative unitarity on the potential’s quartic couplings, i.e., $|\mathrm{Re}\mathcal{M}_{i}|\leqslant\frac{1}{2}$, infers $|e^{j}_{i}|\leqslant 8\pi$.
III.3 Vacuum stability
The vacuum stability requires the potential under the restriction,i.e,$\mathcal{V}>0$ for large field values. The quadratic and cubic terms in the scalar potential can be ignored compared with the quartic term in this limit. In this section, we focus on the vacuum stability induced by dark matter scalars and ignore the mixing quartic couplings with the SM Higgs as well as the triplet scalar for simplicity, and this scenario can work when the mixing quartic couplings with the SM Higgs and triplet scalar take positive or small values as compared to couplings in the dark matter fields. Therefore, the potential with vacuum stability can be given by
$\mathcal{V}_{\mathrm{DM}}=\frac{1}{4}\lambda_{41}\Phi_{1}^{4}+\frac{1}{4}\lambda_{42}\Phi_{2}^{4}+\frac{1}{4}\lambda_{412}\Phi_{1}^{2}\Phi_{2}^{2}+\frac{1}{4}\lambda_{31}\Phi_{1}^{3}\Phi_{2}cos(3\theta_{1}+\theta_{2})+\frac{1}{4}\lambda_{32}\Phi_{1}\Phi_{2}^{3}cos(\theta_{1}-3\theta_{2}).$ (25)
where we have parameterized the dark matter fields with $\phi_{1}=\Phi_{1}e^{i\theta_{1}}$ and $\phi_{2}=\Phi_{2}e^{i\theta_{2}}$.
After minimizing the potential for $3\theta_{1}+\theta_{2}$ and $\theta_{1}-3\theta_{2}$ with $\Phi_{1}\neq 0$ and $\Phi_{2}\neq 0$, the above potential has the formQi and Sun (2023a),
$\mathcal{V}_{\mathrm{DM}}=\frac{1}{4}\lambda_{41}\Phi_{1}^{4}+\frac{1}{4}\lambda_{42}\Phi_{2}^{4}+\frac{1}{4}\lambda_{412}\Phi_{1}^{2}\Phi_{2}^{2}-\frac{1}{4}\lambda_{31}\Phi_{1}^{3}\Phi_{2}-\frac{1}{4}\lambda_{32}\Phi_{1}\Phi_{2}^{3}.$ (26)
Taking $X\equiv\frac{\Phi_{1}}{\Phi_{2}}$, the vacuum stability conditions become
$\lambda_{41}>0,\lambda_{42}>0$ (27)
and
$f(X_{min})>0$ (28)
with
$f(X_{min})=\frac{1}{4}\lambda_{41}X^{4}-\frac{1}{4}\lambda_{31}X^{3}+\frac{1}{4}\lambda_{412}X^{2}-\frac{1}{4}\lambda_{32}X+\frac{1}{4}\lambda_{42},$ (29)
where $X_{min}$ labels the global minimum when $f^{{}^{\prime}}(X_{min})=0$. Then, solving $f^{{}^{\prime}}(X_{min})=0$ by the condition $f^{{}^{\prime\prime}}(X_{min})>0$, we obtain the third condition (28) as
$\lambda_{41}X_{min}^{3}-\frac{3}{4}\lambda_{31}X^{2}+\frac{1}{2}\lambda_{412}X-\frac{1}{4}\lambda_{32}=0,$ (30)
$\lambda_{42}>\frac{1}{4}\lambda_{31}X^{3}_{min}+\frac{1}{2}\lambda_{412}X^{2}_{min}+\frac{3}{4}\lambda_{32}X_{min}.$ (31)
where
$X_{min}=\Bigg{\{}\begin{array}[]{ll}\left(P+\sqrt{P^{2}+Q^{3}}\right)^{1/3}+\left(P-\sqrt{P^{2}+Q^{3}}\right)^{1/3}\,,&D>0\\ 2\sqrt{-Q}\cos\left(\frac{1}{3}\cos^{-1}\left(\frac{P}{\sqrt{-Q^{3}}}\right)\right)\,,&D<0\end{array}$ (34)
with $D\equiv P^{2}+Q^{3}$, $P\equiv\frac{4\lambda_{31}\lambda_{41}\lambda_{412}-8\lambda_{32}\lambda_{41}^{2}-\lambda_{31}^{3}}{64\lambda_{41}^{3}}$, and $Q\equiv\frac{8\lambda_{41}\lambda_{412}-3\lambda^{2}_{31}}{48\lambda_{41}^{2}}$.
IV Dark matter phenomenology
@@ IV.2 Parameter dependence
IV.3 The viable parameter space
@@
The annihilation of dark matter in the late universe is so low that making little difference in the total dark matter abundance. Nevertheless, the effects of such residual annihilations on the cosmic rays can be significant, which enable us detect dark matter indirectly. Indirect detection of dark matter is based on possible signatures of its annihilation such as energetic gamma rays in regions with significant DM density, e.g. the Galactic center, and the strictest bounds come from the Fermi-LATMcDaniel et al. (2024) observation of 30 dSphs for 14.3 years.
V Anti-proton spectrum and electron/positron excess
Figure 12: Background and background + DM (Bkg + DM) of the cosmic-ray antiproton flux with the five benchmark values respectively, where different colored lines correspond to the results of BF taking different values. The data points in both pictures label the PAMELA Adriani et al. (2010) and AMS Aguilar et al. (2019) measurements.
Figure 13: comparison of the positron fraction observed by AMS-02 Aguilar et al. (2014) in the IH scenario (the purple lines) and NH scenario (the blue lines) with the five benchmark values and fixed BF value. The data points stand for the AMS-02, and the green lines represent the Background.
@@ V Anti-proton spectrum and electron/positron excess
Case V $m_{1}=1.1\ \mathrm{TeV},m_{2}=1.21\ \mathrm{TeV},\lambda_{412}=0.01,\mu_{si}=1\ \mathrm{TeV}$, $\lambda_{31}=0.1,\lambda_{32}=0.3,\lambda_{s1h}=0.008,\lambda_{s2h}=0.008,\lambda_{s1d}=0.75,\lambda_{s2d}=1$
@@
Table 2: Benchmark parameters satisfying relic density constraint and direct detection constraint, where Case I (Case III), Case II (Case IV) correspond to the light (heavy) component dominant in dark matter relic density, and Case V is the case that the fraction of $\phi_{1}$ is approximately equal to $\phi_{2}$.
@@
We show the results of the background and Bkg + DM for $\Phi_{\bar{p}}$ in Fig. 12 with the above five cases respectively, where the PAMELA Adriani et al. (2010) results impose strong restrictions on the BF. In Fig. 12(a) and Fig. 12(b), we give the results of the light component dominant in dark matter relic density, while in Fig. 12(c) and Fig. 12(d), we give the heavy component cases. The allowed maximum values of BF are about 2000,400 and 2000 respectively according to pictures (a),(b) and (d) of Fig. 12. For the result of Case III (corresponding to Fig. 12(c)), since the couplings $\lambda_{s1h}$ and $\lambda_{s2h}$ are so small that the contribution of DM particles annihilating to W and Z boson pairs through the s-channel Higgs exchange can be negligible, therefore, the allowed BF value is more flexible compared with other cases, and we have similar results for Case V as in Fig. 12(e), where the fraction of $\phi_{1}$ is approximately equal to $\phi_{2}$ is the total relic density.
@@
In Fig. 13, we give the corresponding results of the comparison of the positron fraction observed by the AMS-02 with the five benchmark values, where the positron fraction is defined by $\Phi_{e^{+}}/(\Phi_{e^{-}}+\Phi_{e^{+}})$. In the model, the final states of DM particle annihilation will depend on whether the neutrino mass is the inverted hierarchy (IH) or the normal hierarchy (NH). The annihilation production is mainly electron final states for the IH scenario, but tau final states for the NH scenario instead. In Fig. 13, the blue lines represent the results of the NH scenario and the purple lines correspond to the IH scenario. When the triplets decay mainly into electron/positron final states, the positron flux arising from DM annihilation will rise sharply above the background as long as we go to higher energies and will eventually drop to the background level, while for the NH scenario, the triplets mostly decay into taus and subsequently decay to electrons, and one can obtain a much softer energy spectrum of the positron flux before eventually dropping to the background level as we can see in Fig. 13(a), Fig. 13(c) and Fig. 13(e). In Fig. 13(b) and Fig. 13(d) corresponding to Case II and Case IV, the results are above the background (without DM) but much lower than the AMS-02 data. The reason is that in those two cases, $m_{1}<M_{\Delta}<m_{2}$, and the annihilation process of $\phi_{1}+\phi_{1}^{\dagger}\to\Delta\Delta$ is kinetically suppressed, while the coupling $\lambda_{s2d}$ related to $\phi_{2}+\phi_{2}^{\dagger}\to\Delta\Delta$ is so small that contribution of dark matter annihilation to triplet is also suppressed. On the other hand, the NH scenario can give a better fit with the AMS-02 compared with the IH scenario as we can see in Fig. 13(a), Fig. 13(c) and Fig. 13(e), which correspond to light component dominant, heavy component dominant and the fraction of $\phi_{1}$ approximately equal to $\phi_{2}$ in the model respectively. As we mentioned above, the positron energy spectrum is harder in the IH scenario since it comes directly from the $\Delta$ decay, while the spectrum is softer in the NH case, because it comes from muon and tau decays. The rise in the AMS-02 positron fraction becomes softer toward higher energies, therefore, it is easier to fit both low- and high-energy bins for the NH case compared to the IH case. In our model, both light components dominant in dark matter relic density and the heavy component case can give the proper results as we can see in Fig. 13(a), Fig. 13(c) and Fig. 13(e) with $M_{\Delta}<m_{1}<m_{2}$, which is consistent with the results of dark matter in the Type-II seesaw mechanism.
VI summary and outlook